# Thermodynamics-Like Formalism for Immiscible and Incompressible Two-Phase Flow in Porous Media

### v1

@@ Abstract
It is possible to formulate immiscible and incompressible two-phase flow in porous media in a mathematical framework resembling thermodynamics based on the Jaynes generalization of statistical mechanics. We review this approach and discuss the meaning of the emergent variables that appear, agiture, flow derivative and flow pressure, which are conjugate to the configurational entropy, the saturation and the porosity respectively. We conjecture that the agiture, the temperature-like variable, is directly related to the pressure gradient. This has as a consequence that the configurational entropy, a measure of how the fluids are distributed within the porous media and the accompanying velocity field, and the mobility of the fluids are related. We also develop elements of another version of the thermodynamics-like formalism where fractional flow rather than saturation is the control variable, since this is typically the natural control variable in experiments.
@@ 1   Introduction
In Section 4, we introduce our approach to immiscible two-phase flow in porous media based on a generalization of statistical mechanics by Jaynes [8, 9]. Statistical mechanics in its traditional role is to connect a molecular description of matter to a description at the much larger scales where matter appears continuous, namely thermodynamics. In adopting this idea to two-phase flow in porous media, we find a thermodynamics-like framework on large scales that together with conservation laws provide a closed set of equations that has the equations of relative permeability theory as a special case. As in ordinary thermodynamics, the difference between intensive and extensive variables is central. We explore this in Section 5. The intensive variables appearing in the thermodynamics-like formalism we derive are emergent in that they have no equivalent on the pore scale. This is discussed in Section 6. The temperature-like intensive variable, agiture, we conjecture to be proportional to the pressure gradient in Subsection 6.1. This conjecture has as a consequence that the configurational entropy, which is an example of Shannon’s information entropy, is related to the mobility of the fluids. The flow derivative we discuss in Subsection 6.2 is the conjugate variable to the saturation which is the fractional volume. We end the section by discussing the conjugate of the porosity, the flow pressure in Subsection 6.3.
@@ 3   Homogenization
It is well known that assuming the three constitutive functions depend only on the saturation is generally not correct. The relative permeabilities do depend not only on the saturation, but also how the fluids arrange themselves in the porous medium, and this depends in turn on the flow rate. The capillary pressure curve is highly hysteretic, signaling that there are missing variables. However, it has turned out to be difficult to go beyond this almost 90 years old phenomenological framework. The dominating approach is based on homogenization. This approach may be divided into two branches: That which focuses on momentum transfer and that which focuses on energy transfer. In the first case, one starts with the hydrodynamic equations on the pore scale [11, 12, 13, 14, 15, 16], and in the second case, one sets up a thermodynamic description at the pore scale [17, 18, 19, 20]. Then comes the coarse-graining, i.e., spatial averaging. There are several ways of doing this, perhaps that of Whitaker is the most well-known [21, 22]. It is based on equating the average of the gradient of a variable associated with pore space to the gradient of the average variable plus an integral over the surface area of the pores. Since the surface area of porous media scales as the volume (i.e., the surface area is extensive — the defining property of porous media), this integral does not vanish as one goes up in scale. We split the variable appearing in the surface integral into an average part and a fluctuating part. This leaves us with having the average, and gradients of the average expressed in terms of the fluctuations of the original variable. The last step is to make an independent assumption on how the fluctuating variables is related to the averages — a closure assumption. With this, the equations between the original, pore-scale variables have been turned into equations between spatial averages of these variables.
@@
The topology of a geometric object such as a porous medium may be described using the four Minkowski functionals volume, surface area, mean curvature, and the Euler characteristics. The Hadwiger theorem states that the Minkowski functionals form a complete basis set for all extensive functions that are invariant with respect to the orientation of the object [28]. The use of this theorem to characterize the free energy of fluids in a porous medium combined with homogenization forms another approach to the scale-up problem [29, 30, 31, 32]. Homogenization as seen so far is a homogenization in space only. McClure et al. [33, 34] emphasize that there is also time which one should average over.
@@ 4   Statistical mechanics
The essence of these efforts has been to find a description of immiscible and incompressible two-phase flow in porous media at scales where the medium acts as a continuum. This is an example of upscaling. The situation is thus that we have a continuum scale phenomenological theory — relative permeability theory that is simple enough to be used in practical situations, but which is approximate. Attempts at going beyond this theory has, however, not led to practical applications due to the complexity that arises. However, if we look beyond the problem we discuss here, there is an example of successful upscaling that is very familiar: The scale-up from a molecular description to thermodynamics based on statistical mechanics. At molecular scales, the position and momentum of each molecule are the variables that describe a monoatomic gas. This works even if we are dealing with millions of molecules in a molecular dynamics simulation [42]. However, it is rather of the order $10^{23}$ molecules that are relevant in e.g. chemistry. A molecular description at such scales is useless. Here emergent variables such as temperature and pressure adequately describe the gas. Statistical mechanics converts the molecular description to thermodynamics which not only contains emergent variables but also relations between them. These variables and their relations have no meaning within a molecular description.
In the mid-fifties, Jaynes generalized statistical mechanics from being a theory for molecular matter to being based on information theory [8]. This was done by replacing Boltzmann’s statistical interpretation of thermodynamic entropy by Shannon’s generalized information entropy [43]. The information entropy is a quantitative measure of what is not known about a system. Suppose we have a stochastic process that can produce $N$ possible outcomes. We number the outcomes from 1 to $N$, assuming a probability $p_{i}$ for the $i$th outcome, $x_{i}$. A quantitative measure of our ignorance about this process must be independent of how we group events together. For example, if we consider composite events such as $p_{ij}=p_{i}p_{j}$, i.e., the probability that an event $i$ is followed by an event $j$, then the information entropy must remain the same. If we know nothing about the process, its information entropy must be at a maximum, since our ignorance is maximal. In order to utilize this, Shannon generalized the Laplace principle of insufficient reason [44] which states that the optimal choice of probabilities for a stochastic process with two possible outcomes is to assign them equal probability, to $N$ outcomes. The optimal choice is to assign all probabilities the same probability, $p_{i}=1/N$. There is only one function that can be constructed from $p_{i}$ with these properties,
@@
Hansen et al. [9] applied the Jaynes generalized statistical mechanics to the immiscible and incompressible two-phase flow in porous media problem. Suppose we have a cylindrical pore sample as shown in Figure 1. Two immiscible fluids enter at the bottom and leaves at the top. The side walls are close off. We assume that the pore structure and chemical composition of the matrix to be statistically uniform throughout the sample. At some distance from the lower edge of the sample, one may assume that the fluids have mixed sufficiently to assume steady-state flow. That is, if we make transversal cuts through the sample, the structure of the porous medium and the distribution of fluids in the pore space to be statistically uniform: Comparing one cut to another, we will not be able to determine which is closest to the inlet. Figure 1 shows one such cut. Within this cut, we pick out a Representative Elementary Area (REA), which is large enough to represent the statistics of the sample and small enough to have the flow fluctuate. There is an volumetric flow rate $Q_{p}$ passing through the REA which may be split into a volumetric flow rate of the wetting fluid $Q_{w}$ and the volumetric flow rate of the non-wetting fluid $Q_{n}$, i.e.,
@@
At a given time for a given REA, the fluid configuration, characterized by the velocity field and where the fluids are, is $X$. The quantities that describe the REA depend on $X$ so that we have $Q_{p}(X)$, $Q_{w}(X)$, $Q_{n}(X)=Q_{p}(X)-Q_{w}(X)$, $A_{p}(X)$, $A_{w}(X)$, and $A_{n}(X)=A_{p}(X)-A_{n}(X)$. Of course, $A_{p}(X)$ is not dependent on the velocity field, nor the distribution of the two fluids, but only on the shape of pore space. Let us denote $p(X)$ the probability density to find fluid and pore space configuration $X$. We define a configurational entropy [^1] The probability density $p(X)$ is continuous, so that we are dealing with differential entropy [45] in equation (13). Another way of dealing with the transition from discrete to continuous systems is that of Jaynes [46].
@@
We may now calculate the entropy $S$ by inserting equation (18) in equation (13) and using equation (19). We obtain
@@
$Q_{z}=Q_{u}-\frac{1}{\lambda_{u}}\ S+\frac{\lambda_{w}}{\lambda_{u}}\ A_{w}++\frac{\lambda_{p}}{\lambda_{u}}\ A_{p}\;.$ (21)
@@
where we have used that this is a Legendre transform since
@@ 6.1   Agiture
Figure 3 A cylindrical porous medium consisting of two halves A and B. They have different properties such as porosity, chemical composition of the matrix, or the pore spaces have different topology. At the bottom one third of the porous cylinder, the boundary between the two halves is impenetrable. This is indicated by the dark section of the boundary. Above, the two halves are in direct contact. We inject the immiscible fluids into each of the halves A and B at the bottom. We may do this by injecting at different volumetric flow rates $Q_{p}^{A}=Q_{w}^{A}+Q_{n}^{A}$ and $Q_{p}^{B}=Q_{w}^{B}+Q_{n}^{B}$. This gives rise to flow derivatives $\mu_{A}$ and $\mu_{B}$, and agitures $\theta_{A}$ and $\theta_{B}$ in the two halves far enough from the inlet for the flow to be in a steady state. The pressure gradient will then point along the axis of the cylinder, the $z$-direction, and we will in the following use the notation $P^{\prime}=\partial P/\partial z$ for the gradient. We find a pressure gradient $P^{\prime}_{A}$ and $P^{\prime}_{B}$ in A and B respectively. Further into the cylindrical sample above the impenetrable wall, the flow will adjust to a new steady state characterized by equal agiture $\theta$ and flow derivative $\mu$ and in both halves according to equations (38) and (39). Likewise the pressure gradient $P^{\prime}$ must be equal in the two halves since there cannot be any net flow in either direction across the boundary between the two halves, apart from local fluctuations.
@@
We note that the mobility is defined as
@@
Using the relative permeability constitutive equation (43), we find the mobility in this case to be
@@
Comparing equations (46) and (47), we see that [^2] Since we are dealing with differential entropy, it is not a problem that it is negative. Suppose we have that $\mu_{w}=\mu_{n}$. The mobility will then have a minimum for the saturation that leads to the most fluid-fluid interfaces. It makes sense that this is also the saturation for which the entropy is at a maximum.
@@
It is surprising that the mobility and the entropy should be related.
@@ 6.3   Flow pressure
$\mu=-\left(\frac{\partial Q_{p}}{\partial A_{p}}\right)_{\theta,A_{w}}=\left(\frac{\partial v_{p}}{\partial\phi}\right)_{\theta,S_{w}}\;.$ (56)
@@ 7   Seepage velocities and the co-moving velocity
We note that the thermodynamic velocities $({\hat{v}}_{w},{\hat{v}}_{n})$ can be found by knowing $v_{p}$ alone, see equations (52) and (53). The seepage velocities, on the other hand requires additional knowledge, i.e., $v_{m}$, and are given by equations (69) and (70). They provide a two-way mapping
@@ 9   Discussion and conclusion
We have conjectured here that the agiture, the temperature-like emergent variable conjugate to the configurational entropy, is proportional to the pressure gradient. A surprising consequence of this conjecture is that the mobility of the fluids is directly related to the entropy: higher mobility means lower configurational entropy.

### v2

@@ Abstract
It is possible to formulate immiscible and incompressible two-phase flow in porous media in a mathematical framework resembling thermodynamics based on the Jaynes generalization of statistical mechanics. We review this approach and discuss the meaning of the emergent variables that appear, agiture, flow derivative and flow pressure, which are conjugate to the configurational entropy, the saturation and the porosity respectively. We conjecture that the agiture, the temperature-like variable, is directly related to the pressure gradient. This has as a consequence that the configurational entropy, a measure of how the fluids are distributed within the porous media and the accompanying velocity field, and the differential mobility of the fluids are related. We also develop elements of another version of the thermodynamics-like formalism where fractional flow rather than saturation is the control variable, since this is typically the natural control variable in experiments.
@@ 1   Introduction
In Section 4, we introduce our approach to immiscible two-phase flow in porous media based on a generalization of statistical mechanics by Jaynes [8, 9]. Statistical mechanics in its traditional role is to connect a molecular description of matter to a description at the much larger scales where matter appears continuous, namely thermodynamics. In adopting this idea to two-phase flow in porous media, we find a thermodynamics-like framework on large scales that together with conservation laws provide a closed set of equations that has the equations of relative permeability theory as a special case. As in ordinary thermodynamics, the difference between intensive and extensive variables is central. We explore this in Section 5. The intensive variables appearing in the thermodynamics-like formalism we derive are emergent in that they have no equivalent on the pore scale. This is discussed in Section 6. The temperature-like intensive variable, agiture, we conjecture to be proportional to the pressure gradient in Subsection 6.1. This conjecture has as a consequence that the configurational entropy, which is an example of Shannon’s information entropy, is related to the differential mobility [^1] The differential mobility is the derivative of the flow velocity with respect to the pressure gradient. The mobility is the flow velocity divided by the pressure gradient. of the fluids. The flow derivative we discuss in Subsection 6.2 is the conjugate variable to the saturation which is the fractional volume. We end the section by discussing the conjugate of the porosity, the flow pressure in Subsection 6.3.
@@ 3   Homogenization
It is well known that assuming the three constitutive functions depend only on the saturation is generally not correct. The relative permeabilities do depend not only on the saturation, but also how the fluids arrange themselves in the porous medium, and this depends in turn on the flow rate. The capillary pressure curve is highly hysteretic, signaling that there are missing variables. However, it has turned out to be difficult to go beyond this almost 90 years old phenomenological framework. The dominating approach is based on homogenization. This approach may be divided into two branches: That which focuses on momentum transfer and that which focuses on energy transfer. In the first case, one starts with the hydrodynamic equations on the pore scale [11, 12, 13, 14, 15, 16], and in the second case, one sets up a thermodynamic description at the pore scale [17, 18, 19, 20]. Then comes the coarse-graining, i.e., spatial averaging. There are several ways of doing this, perhaps that of Whitaker is the most well-known [21, 22]. It is based on equating the average of the gradient of a variable associated with pore space to the gradient of the average variable plus an integral over the surface area of the pores. Since the surface area of porous media scales as the volume (i.e., the surface area is extensive — the defining property of porous media), this integral does not vanish as one goes up in scale. We split the variable appearing in the surface integral into an average part and a fluctuating part. This leaves us with having the average, and gradients of the average expressed in terms of the fluctuations of the original variable. The last step is to make an independent assumption on how the fluctuating variables is related to the averages –– a closure assumption. With this, the equations between the original, pore-scale variables have been turned into equations between spatial averages of these variables.
@@
The topology of a geometric object such as a porous medium may be described using the four Minkowski functionals volume, surface area, mean curvature, and the Euler characteristics. The Hadwiger theorem states that the Minkowski functionals form a complete basis set for all extensive functions that are invariant with respect to the orientation of the object [28]. The use of this theorem to characterize the free energy of fluids in a porous medium combined with homogenization forms another approach to the scale-up problem [29, 30, 31, 32]. Homogenization as seen so far is based on spatial averages only. McClure et al. [33, 34] emphasize that there is also time which one should average over and that different processes work on different time scales.
@@ 4   Statistical mechanics
The essence of these efforts has been to find a description of immiscible and incompressible two-phase flow in porous media at scales where the medium acts as a continuum. [^2] Including compressibility would necessitate switching from conserved fluid volume to conserved fluid mass. It would also couple the theory to ordinary thermodynamics through equations of state for the compressible fluids. We wish to avoid these complications at the present stage. This is an example of upscaling. The situation is thus that we have a continuum scale phenomenological theory — relative permeability theory that is simple enough to be used in practical situations, but which is approximate. Attempts at going beyond this theory has, however, not led to practical applications due to the complexity that arises. However, if we look beyond the problem we discuss here, there is an example of successful upscaling that is very familiar: The scale-up from a molecular description to thermodynamics based on statistical mechanics. At molecular scales, the position and momentum of each molecule are the variables that describe a monoatomic gas. This works even if we are dealing with millions of molecules in a molecular dynamics simulation [42]. However, it is rather of the order $10^{23}$ molecules that are relevant in e.g. chemistry. A molecular description at such scales is useless. Here emergent variables such as temperature and pressure adequately describe the gas. Statistical mechanics converts the molecular description to thermodynamics which not only contains emergent variables but also relations between them. These variables and their relations have no meaning within a molecular description.
In the mid-fifties, Jaynes generalized statistical mechanics from being a theory for molecular matter to being based on information theory [8]. This was done by replacing Boltzmann’s statistical interpretation of thermodynamic entropy by Shannon’s generalized information entropy [43]. The information entropy is a quantitative measure of what is not known about a system. Suppose we have a stochastic process that can produce $N$ possible outcomes. We number the outcomes from 1 to $N$, assuming a probability $p_{i}$ for the $i$th outcome, $x_{i}$. A quantitative measure of our ignorance about this process must be independent of how we group events together. For example, if we consider composite events such as $p_{ij}=p_{i}p_{j}$, i.e., the probability that an event $i$ is followed by an event $j$, then the information entropy must remain the same. If we know nothing about the process, its information entropy must be at a maximum, since our ignorance is maximal. In order to utilize this, Shannon generalized to $N$ outcomes the Laplace principle of insufficient reason [44] which states that the optimal choice of probabilities for a stochastic process with two possible outcomes is to assign them equal probability. Hence, the optimal choice when there are $N$ outcomes is to assign all probabilities the same value, $p_{i}=1/N$. There is only one function that can be constructed from $p_{i}$ with these properties,
@@
Jaynes provided a set of circumstances for his generalized statistical mechanics to be applicable. Hansen et al. [9] demonstrated that these circumstances are met for steady-state immiscible two-phase flow in porous media, and hence proceeded to construct such a statistical mechanics. The generalized statistical mechanics will produce a thermodynamics-like framework at large scales. This thermodynamic-like framework is the end goal of this approach.
Suppose we have a cylindrical pore sample as shown in Figure 1. Two immiscible fluids enter at the bottom and leaves at the top. The side walls are close off. We assume that the pore structure and chemical composition of the matrix to be statistically uniform throughout the sample. At some distance from the lower edge of the sample, one may assume that the fluids have mixed sufficiently to assume steady-state flow. That is, if we make transversal cuts through the sample, the structure of the porous medium and the distribution of fluids in the pore space to be statistically uniform: Comparing one cut to another, we will not be able to determine which is closest to the inlet. Figure 1 shows one such cut. Within this cut, we pick out a Representative Elementary Area (REA), which is large enough to represent the statistics of the sample and small enough to have the flow fluctuate. There is an volumetric flow rate $Q_{p}$ passing through the REA which may be split into a volumetric flow rate of the wetting fluid $Q_{w}$ and the volumetric flow rate of the non-wetting fluid $Q_{n}$, i.e.,
@@
At a given time for a given REA, the fluid configuration, characterized by the velocity field and where the fluids are, is $X$. The quantities that describe the REA depend on $X$ so that we have $Q_{p}(X)$, $Q_{w}(X)$, $Q_{n}(X)=Q_{p}(X)-Q_{w}(X)$, $A_{p}(X)$, $A_{w}(X)$, and $A_{n}(X)=A_{p}(X)-A_{w}(X)$. Of course, $A_{p}(X)$ is not dependent on the velocity field, nor the distribution of the two fluids, but only on the shape of pore space. Let us denote $p(X)$ the probability density to find fluid and pore space configuration $X$. We define a configurational entropy [^3] The probability density $p(X)$ is continuous, so that we are dealing with differential entropy [46] in equation (13). Another way of dealing with the transition from discrete to continuous systems is that of Jaynes [47].
@@
The partition function is in some sense the “end product" of a statistical mechanics approach to a problem. It is the generating function for the probability distribution for microscopic configurations. With it, all macroscopic averages may be calculated.
In ordinary statistical mechanics, a hamiltonian, i.e., a microscopic energy function is necessary input for calculating the partition function. In the present problem we have three pore-level functions, $Q_{u}(X)$, $A_{w}(X)$ and $A_{p}(X)$. The problem in ordinary statistical mechanics, is that even for seemingly simple problems such as the two-dimensional Ising model for magnetic systems, calculating the partition function may be extremely difficult, cfr. Onsager’s calculation of the partition function for this system [45]. The problem is surely no less difficult in the porous media context. Perhaps, it is even worse in that we do not have good models for the three microscopic functions in contrast to ordinary statistical mechanics where model hamiltonians are legio.
A byproduct of the statistical mechanics approach is the emergence of a set of thermodynamics-like relations between the REA-scale variables. This framework will be in place whether it is possible to calculate the partition function or not. We will now develop this thermodynamics-like framework for immiscible two-phase flow in porous media.
We start by calculating the entropy $S$ by inserting equation (17) in equation (13) and using equation (19). We obtain
@@
$Q_{z}=Q_{u}-\frac{1}{\lambda_{u}}\ S+\frac{\lambda_{w}}{\lambda_{u}}\ A_{w}+\frac{\lambda_{p}}{\lambda_{u}}\ A_{p}\;.$ (21)
@@
where we have used that this is a Legendre transform [^4] Assuming convexity of the involved functions. since
@@ 6.1   Agiture
Figure 3 shows a cylindrical porous medium consisting of two halves A and B. They have different properties such as porosity, chemical composition of the matrix, or the pore spaces have different topology. At the bottom one third of the porous cylinder, the boundary between the two halves is impenetrable. This is indicated by the dark section of the boundary. Above, the two halves are in direct contact. We inject the immiscible fluids into each of the halves A and B at the bottom. We may do this by injecting at different volumetric flow rates $Q_{p}^{A}=Q_{w}^{A}+Q_{n}^{A}$ and $Q_{p}^{B}=Q_{w}^{B}+Q_{n}^{B}$. This gives rise to flow derivatives $\mu_{A}$ and $\mu_{B}$, and agitures $\theta_{A}$ and $\theta_{B}$ in the two halves far enough from the inlet for the flow to be in a steady state. The pressure gradient will then point along the axis of the cylinder, the $z$-direction, and we will in the following use the notation $P^{\prime}=\partial P/\partial z$ for the gradient. We find a pressure gradient $P^{\prime}_{A}$ and $P^{\prime}_{B}$ in A and B respectively. Further into the cylindrical sample above the impenetrable wall, the flow will adjust to a new steady state characterized by equal agiture $\theta$ and flow derivative $\mu$ and in both halves according to equations (38) and (39). Likewise the pressure gradient $P^{\prime}$ must be equal in the two halves since there cannot be any net flow in either direction across the boundary between the two halves, apart from local fluctuations.
@@
We note that the differential mobility is defined as
@@
Using the relative permeability constitutive equation (43), we find the mobility [^5] Mobility and differential mobility are equal in this case as the velocity is linear in the pressure gradient. in this case to be
@@
Comparing equations (46) and (47), we see that [^6] Since we are dealing with differential entropy, it is not a problem that it is negative.
@@
It is surprising that the differential mobility and the entropy should be related.
Equation (41) relates the agiture $\theta$ and the pressure gradient $P^{\prime}$ and the way we have argued is by demonstrating that the pressure gradient behaves as one would expect an agiture (= agitation temperature) should do. However, could one test equation (41) directly by measuring the left hand side and compare it to the right hand side. This is difficult as it would entail measuring the configurational entropy and then use equation (44); and measuring entropy is difficult, also in ordinary thermodynamics.
@@ 6.3   Flow pressure
$\pi=-\left(\frac{\partial Q_{p}}{\partial A_{p}}\right)_{\theta,A_{w}}=\left(\frac{\partial v_{p}}{\partial\phi}\right)_{\theta,S_{w}}\;.$ (56)
@@
We expect the flow pressure $\pi$ to be of less practical use, as we expect that the flow rate $Q_{p}$ to be proportional to the pore area $A_{p}$. This makes the Legendre transform between $A_{p}$ and $\pi$ break down due to lack of convexity. The flow pressure will nevertheless play a role e.g., when porosity gradients are present.
@@ 7   Seepage velocities and the co-moving velocity
We note that the thermodynamic velocities $({\hat{v}}_{w},{\hat{v}}_{n})$ can be found by knowing $v_{p}$ alone, see equations (52) and (53). The seepage velocities, on the other hand requires additional knowledge, i.e., $v_{p}$ and $v_{m}$, and are given by equations (69) and (70). They provide a two-way mapping
@@ 9   Discussion and conclusion
We noted in Section 4 that actually calculating the partition function, equation (19), may be impossible in the present context. The consequence of this is that we are unable to derive directly the constitutive equation $Q_{p}(P^{\prime},A_{w},A_{p})$ from the pore level physics. It must be found by other means, and then used as input in the thermodynamics-like formalism.
We have conjectured here that the agiture, the temperature-like emergent variable conjugate to the configurational entropy, is proportional to the pressure gradient. A surprising consequence of this conjecture is that the differential mobility of the fluids is directly related to the entropy: higher differential mobility means lower configurational entropy.
@@
We thank Steffen Berg for valuable discussions and comments.